Find the maximum possible area of a rectangle in quadrant 1 under the curve y = (x − 6)^2. (Include
a test showing that your rectangle’s area is the maximum possible.)
Answer #1

Please see the solution and if you have any problem please comment

Thanks

Solution y = (x-6)² y (0,36) Area of rectangle = (x) (y) Ary (6,0) y = (x-6)² A = x(x-6)2 = x (x² x 36-12x) A = x +36*-12x² f